# MSC1, MSC2, MSC3

### v1

@@ 4 Numerical examples
    In the first case we consider a small dimensional problem with $d=400$, so we have a system with 801 masses. Here we damp the 9 lowest undamped eigenfrequencies, i.e., $s=9$ in (30). The damping geometry is determined by (38) and 20 different damping positions will be considered. These are determined by the following indices:
@@
    In the second case we increase the problem dimension by considering $d=1000$, thus here we have a system with 2001 masses and $n=4002$. In this case, we damp the 21 lowest undamped eigenfrequencies, i.e., $s=21$ in (30). Here we consider 25 damping positions determined by the following indices:
@@
    We apply the projection framework illustrated in Section 2.2 for the solution of this larger problem. We would like to mention that a Lyapunov equation whose coefficient matrix is of dimension $4002$ as in this case is usually not considered to be large-scale in the matrix equation community. On the other hand, the huge number of Lyapunov equations we need to solve within the optimization procedure makes our novel projection framework very appealing. In Algorithm 1 we employ $\epsilon=10^{-8}$ and $m_{\max}=120$. The projected equations (19) are solved by means of our recycling Krylov approach adopting the same parameters as in case a) above.
@@
We first focus on case a) described above and in Figure 2 we report the relative errors obtained by our recycling Krylov method. The relative errors in the optimal viscosities were calculated by $\|\mathsf{v}^{*}-\mathsf{v}\|/\|\mathsf{v}\|$, where $\mathsf{v}$ and $\mathsf{v}^{*}$ denote the optimal viscosity vectors calculated by the recycling Krylov method and the Matlab’s function lyap, respectively. Similarly, the relative errors in the average total energy are calculated by $|\operatorname*{trace}({X(\mathsf{v})}^{*})-\operatorname*{trace}(X(\mathsf{v}))|/\operatorname*{trace}({X(\mathsf{v})}^{*})$, where $\operatorname*{trace}({X(\mathsf{v})}^{*})$ is once again the optimal trace for the given configuration obtained by the Matlab’s function lyap, and $\operatorname*{trace}(X(\mathsf{v}))$ is the optimal trace calculated by our recycling Krylov scheme.
@@
Figure 2: Example 4.1, the case a). Relative errors in the total average energy (squares) and in the viscosity (circles) at optimal gains for the recycling Krylov method.
@@
From the results reported in Figure 2 we can notice that our novel recycling Krylov approach leads to a solution process able to achieve very small errors in the average total energy. Also the computed optimal viscosities turn out to be rather accurate attaining an error of the same order of the threshold used for fminsearch. Moreover, our novel recycling Krylov strategy is very efficient. In particular, in this example, it accelerated the overall optimization process approximately 2.7 times (in the average case). Figure 3 shows a precise acceleration ratio for each considered configuration. The displayed time ratio is the ratio between the total time needed for the direct calculation of the average total energy by using Matlab’s function lyap and the total time needed for the viscosity optimization by using the recycling Krylov method.
@@
We would like to mention that, for this example, we tried to use also GCRO-DR with no preconditioning within our recycling Krylov technique. In the best case scenarios, plain GCRO-DR managed to converge by performing a larger number of iterations: $\mathcal{O}(10^{2})$ iterations to be compared to the $\mathcal{O}(10^{1})$ iterations performed in the preconditioned case. Such large number of iterations remarkably worsened the overall performance in terms of computational time. In some other cases, unpreconditioned GCRO-DR did not achieve the prescribed level of accuracy thus jeopardizing the convergence of the outer minimization procedure. This shows that the preconditioning operator $\mathcal{P}$ described in (15) works well even though more performing preconditioners can certainly be designed.
@@
In the next numerical examples we are going to illustrate the following two cases of analysis of multi-agent systems that are of interest:
@@
  • We choose $H=I_{m}$. For a given set of subsets of agents $\left\{J_{1},\ldots,J_{q}\right\}$, $J_{i}\subset\left\{1,\ldots,m\right\}$, and a given set of parameters of agents in $J_{i}$, we study the influence of these parameters to the dynamics of the overall system. By calculating the corresponding $H_{2}$-norms we can study the influence of these parameters on the dynamics of the whole system when all agents are exposed to independent external disturbances.
@@
$A_{i}=\begin{bmatrix}-10&5\\ 5&-8\end{bmatrix},\quad\text{for}\quad i=1,\ldots,m.$
@@
$K_{i}=\begin{bmatrix}0.3&0\\ 0&0.2\end{bmatrix},\quad B_{i}=C_{i}^{\mkern-1.5mu\mathsf{T}}=\begin{bmatrix}1&1\\ -1&1\end{bmatrix},\quad\mbox{for}\quad i=1,\ldots,m.$
@@
$\mathbf{A}_{ij}$ $=\delta_{i,j}A_{i}-\ell_{ij}B_{i}K_{i}C_{j},\quad i,j\in\left\{1,\ldots,m\right\}.$
@@
$\mathbf{A}^{(k)}(\mathsf{v})=\mathbf{A}-B_{l}\mathop{\mathrm{diag}}(v_{1},v_{2},v_{2},v_{3})B_{r},$ (39)
where $B_{l},B_{r}\in{\mathbb{R}}^{m\times 4}$ are determined by the agent index $k$. It holds
@@
whereas all other entries of $B_{l}$ and $B_{r}$ are zero. In this example we considered the influence of the variation of the dynamics of one agent by calculating the $H_{2}$-norm of a corresponding system when only this agent is externally disturbed. Thus, in the case of the matrix $\mathbf{A}^{(k)}$ (that corresponds to analyzing the $k$-th agent), the corresponding matrix $\mathbf{E}$ is given by $\mathbf{E}=e_{k}\otimes I_{n}$, for $k=1,\ldots,m$.
@@
We note that some of the instances of $\mathbf{A}^{(k)}(\mathsf{v})$ turned out to be non stable. We have discarded those instances in our analysis. By doing so, for this particular example, the total number of considered parameters is equal to 1 400 328. This means that we needed to calculate the $H_{2}$-norm, and thus solve a Lyapunov equation, a significant number of times. In particular, for fixed $v_{1}$, $v_{2}$, $v_{3}$, we are interested in identifying for which $k$, $k=1,\ldots,m$, the $H_{2}$-norm of the underlying system over described parameter variations is maximal. This provides an insight on the importance of the considered agent.
Table 2 presents the average computational time needed by different methods to perform the task described above. We report the average required time by our novel solvers, namely the recycling Krylov method and the projection framework (denoted in the table by rec. Krylov and proj. framework, respectively). The computational parameters of our recycling Krylov method are as in Example 4.1. The only exceptions are in the GCRO-DR threshold and the rank $\bar{p}$ of the TSVD $\mathcal{N}_{1}\mathcal{M}_{1}^{\mkern-1.5mu\mathsf{T}}$ and $\mathcal{N}_{2}\mathcal{M}_{2}^{\mkern-1.5mu\mathsf{T}}$ used in the preconditioner. Here we use $10^{-8}$ and $\bar{p}=5$, respectively. For Algorithm 1 we used $\epsilon=10^{-10}$ and $m_{\max}=200$. The projected equations are solved by our recycling Krylov method with the same setting we have just described above.
@@
Figure 6 illustrates the topology of the problem and provides information regarding the maximal $H_{2}$-norm we computed by varying the parameter of the system. This latter aspect is illustrated by the node colors. In particular, darker colors correspond to the agents where the maximal $H_{2}$-norm turns out to be larger. It can be seen that certain groups of agents result in much larger $H_{2}$-norms. Therefore, they are much more important for the considered multi-agent system as altering those agents may lead to an (almost) unstable system. Thanks to our novel solvers, this kind of analysis can now be accurately carried out at acceptable cost.
@@
In this example, we still consider agent dynamics defined by $2\times 2$ state matrices of dimension $m$. We aim at analyzing the parameter variation in off-diagonal elements of consecutive agents. The system matrices $A_{i}$, $B_{i}$, $C_{i}$, $K_{i}$, $C_{i}$, and $\mathbf{A}$ are defined as in Example 4.2 but here we consider $m=200$.
The Laplacian matrix $L$ of the underlying graph is given by Algorithm 2 and the corresponding matrices $B_{l}$ and $B_{r}$ determining the low-rank perturbation are defined in the following way. For a given odd index $31\leqslant k\leqslant 2m-$ 3, $B_{l},B_{r}\in{\mathbb{R}}^{m\times 4}$ are
@@ 10       $L(i,i)=-\sum_{j=1}^{m}(L(i,j))$;
$\mathbf{A}^{(k)}(\mathsf{v})=\mathbf{A}-B_{l}\mathop{\mathrm{diag}}(v_{1},v_{1},v_{2},v_{2})B_{r}.$
@@
We will study the behavior of the $H_{2}$-norm of the underlying system by varying the parameter $k\in\{61,141,221,301\}$. The total number of parameter configurations $\mathsf{v}=(v_{1},v_{2})$ for which $\mathbf{A}^{k}(\mathsf{v})$ is stable is equal to 6 04460446\,0446 044. Moreover, in this example, we considered the matrix $\mathbf{E}=I_{2m}$ for all cases, meaning that we measured the $H_{2}$-norm of the corresponding systems when all agents were independently externally disturbed.
@@
From the results in Table 3 we can see that our recycling Krylov scheme always attains very small relative errors. On the other hand, it turns out to be not very competitive in terms of computational timings. A finer tuning of the GCRO-DR parameters and the preconditioning operator may lead to some improvements in the performance of our scheme. Our projection framework outperforms all other approaches in terms of computational timing.
@@
In Figure 7 we display the magnitude of the $H_{2}$-norm by varying $\mathsf{v}$ for a fixed $k$. It can be seen how certain parameter values have a much greater impact on the $H_{2}$-norm of the system thus emphasizing the important role of the corresponding pairs of agents. Once again, thanks to our novel solution processes, such analysis can be carried out in very few seconds on a standard laptop.
@@ 5 Conclusions
We proposed two different, efficient, and accurate methods for solving sequences of parametrized Lyapunov equations. The recycling Krylov approach is well-suited for small dimensional problems and is able to provide solutions achieving a very small relative errors. The proposed projection framework relies on the extended Krylov subspace method, and it makes use of the aforementioned recycling Krylov technique to solve the projected equations. Even though this second algorithm is tailored to large-scale problems, our numerical results show that it is extremely competitive also for medium-sized problems, especially if the number of Lyapunov equations to be solved is very large as it happens in the application settings we studied. We showed that the projection framework is able to speed up the entire solution process by an order of magnitude. Expensive analyses like viscosity optimization for vibrational systems and output synchronization of multi-agent systems are, thus, now affordable.
@@ Acknowledgements
The first author is member of the INdAM Research Group GNCS. His work was partially supported by the funded project GNCS2023 “Metodi avanzati per la risoluzione di PDEs su griglie strutturate, e non” (CUP_E53C22001930001).

### v2

@@ 4 Numerical examples
    In the first case, we consider a small dimensional problem with $d=400$, so we have a system with 801 masses. Here we damp the 9 lowest undamped eigenfrequencies, i.e., $s=9$ in (29). The damping geometry is determined by (37) and 20 different damping positions will be considered. These are determined by the following indices:
@@
    In the second case, we increase the problem dimension by considering $d=1000$, thus here we have a system with 2001 masses and $n=4002$. In this case, we damp the 21 lowest undamped eigenfrequencies, i.e., $s=21$ in (29). Here we consider 25 damping positions determined by the following indices:
@@
    We apply the projection framework illustrated in Section 2.2 for the solution of this larger problem. We would like to mention that a Lyapunov equation whose coefficient matrix is of dimension $4002$ as in this case is usually not considered to be large scale in the matrix equation community. On the other hand, the huge number of Lyapunov equations we need to solve within the optimization procedure makes our novel projection framework very appealing. In Algorithm 1 we employ $\epsilon=10^{-8}$ and $m_{\max}=120$. The projected equations (18) are solved by means of our SMW+recycling Krylov approach, adopting the same parameters as in case a) above.
@@
We first focus on case a) described above and in Figure 2 we report the relative errors obtained by our SMW+recycling Krylov method. The relative errors in the optimal viscosities were calculated by $\|\mathsf{v}^{*}-\mathsf{v}\|/\|\mathsf{v}\|$, where $\mathsf{v}$ and $\mathsf{v}^{*}$ denote the optimal viscosity vectors calculated by the SMW+recycling Krylov method and the Matlab’s function lyap, respectively. Similarly, the relative errors in the average total energy are calculated by $|\operatorname*{trace}({X(\mathsf{v})}^{*})-\operatorname*{trace}(X(\mathsf{v}))|/\operatorname*{trace}({X(\mathsf{v})}^{*})$, where $\operatorname*{trace}({X(\mathsf{v})}^{*})$ is once again the optimal trace for the given configuration obtained by the Matlab’s function lyap, and $\operatorname*{trace}(X(\mathsf{v}))$ is the optimal trace calculated by our SMW+recycling Krylov scheme.
@@
Figure 2: Example 4.1, the case a). Relative errors in the total average energy (squares) and in the viscosity (circles) at optimal gains for the SMW+recycling Krylov method.
@@
From the results reported in Figure 2 we can notice that our novel SMW+recycling Krylov approach leads to a solution process able to achieve very small errors in the average total energy. Also, the computed optimal viscosities turn out to be rather accurate, attaining an error of the same order of the threshold used for fminsearch. Moreover, our novel SMW+recycling Krylov strategy is very efficient. In particular, in this example, it accelerated the overall optimization process approximately 2.7 times (in the average case). Figure 3 shows a precise acceleration ratio for each considered configuration. The displayed time ratio is the ratio between the total time needed for the direct calculation of the average total energy by using Matlab’s function lyap and the total time needed for the viscosity optimization by using the SMW+recycling Krylov method.
@@
We would like to mention that, for this example, we tried to use also GCRO-DR with no preconditioning within our recycling Krylov technique. In the best case scenarios, plain GCRO-DR managed to converge by performing a larger number of iterations: $\mathcal{O}(10^{2})$ iterations to be compared to the $\mathcal{O}(10^{1})$ iterations performed in the preconditioned case. Such large number of iterations remarkably worsened the overall performance in terms of computational time. In some other cases, unpreconditioned GCRO-DR did not achieve the prescribed level of accuracy, thus jeopardizing the convergence of the outer minimization procedure. This shows that the preconditioning operator $\mathcal{P}$ described in (15) works well, even though more performing preconditioners can certainly be designed.
@@
In the next numerical examples, we are going to illustrate the following two cases of analysis of multi-agent systems that are of interest:
@@
  • We choose $H=I_{m}$. For a given set of subsets of agents $\left\{J_{1},\ldots,J_{q}\right\}$, $J_{i}\subset\left\{1,\ldots,m\right\}$, and a given set of parameters of agents in $J_{i}$, we study the influence of these parameters on the dynamics of the overall system. By calculating the corresponding $H_{2}$-norms, we can study the influence of these parameters on the dynamics of the whole system when all agents are exposed to independent external disturbances.
@@
$A^{(i)}=\begin{bmatrix}-10&5\\ 5&-8\end{bmatrix},\quad\text{for}\quad i=1,\ldots,m.$
@@
$K^{(i)}=\begin{bmatrix}0.3&0\\ 0&0.2\end{bmatrix},\quad B^{(i)}=(C^{(i)})^{\mkern-1.5mu\mathsf{T}}=\begin{bmatrix}1&1\\ -1&1\end{bmatrix},\quad\mbox{for}\quad i=1,\ldots,m.$
@@
$\mathbf{A}_{ij}$ $=\delta_{i,j}A^{(i)}-\ell_{ij}B^{(i)}K^{(i)}C^{(j)},\quad i,j\in\left\{1,\ldots,m\right\}.$
@@
$\mathbf{A}^{(k)}(\mathsf{v})=\mathbf{A}-B_{l}\mathop{\mathrm{diag}}(v_{1},v_{2},v_{2},v_{3})B_{r}^{\mkern-1.5mu\mathsf{T}},$ (38)
where $B_{l},B_{r}\in{\mathbb{R}}^{2m\times 4}$ are determined by the agent index $k$. It holds,
@@
whereas all other entries of $B_{l}$ and $B_{r}$ are zero. In this example, we considered the influence of the variation of the dynamics of one agent by calculating the $H_{2}$-norm of a corresponding system when only this agent is externally disturbed. Thus, in the case of the matrix $\mathbf{A}^{(k)}$ (that corresponds to analyzing the $k$-th agent), the corresponding matrix $\mathbf{E}$ is given by $\mathbf{E}=e_{k}\otimes I_{n}$, for $k=1,\ldots,m$.
@@
We note that some instances of $\mathbf{A}^{(k)}(\mathsf{v})$ turned out to be non-stable. In particular, we have computed the rightmost eigenvalue $\lambda$ of $\mathbf{A}^{(k)}(\mathsf{v})$ and discarded those configurations for which $\text{Re}(\lambda)\geqslant 0$. By doing so, for this particular example, the total number of considered parameters is equal to 1 400 328444This number is obtained by multiplying all the possible values of $v_{1}$, $v_{2}$, and $v_{3}$, namely $21^{3}$, by all the adopted selections of $k$, i.e., $m$, and then subtracting the number of unstable cases (414 828414828414\,828414 828).. This means that we needed to calculate the $H_{2}$-norm, and thus solve a Lyapunov equation, a significant number of times. In particular, for fixed $v_{1}$, $v_{2}$, $v_{3}$, we are interested in identifying for which $k$, $k=1,\ldots,m$, the $H_{2}$-norm of the underlying system over described parameter variations is maximal. This provides an insight on the importance of the considered agent.
Table 2 presents the average computational time needed by different methods to perform the task described above. We report the average required time by our novel solvers, namely the SMW+recycling Krylov method and the projection framework (denoted in the table by SMW+rec. Krylov and proj. framework, respectively). The computational parameters of our SMW+recycling Krylov method are as in Example 4.1. The only exceptions are in the GCRO-DR threshold and the rank $\bar{p}$ of the TSVD $\mathcal{N}_{1}\mathcal{M}_{1}^{\mkern-1.5mu\mathsf{T}}$ and $\mathcal{N}_{2}\mathcal{M}_{2}^{\mkern-1.5mu\mathsf{T}}$ used in the preconditioner. Here, we use $10^{-8}$ and $\bar{p}=5$, respectively. For Algorithm 1 we used $\epsilon=10^{-10}$ and $m_{\max}=200$. The projected equations are solved by our SMW+recycling Krylov method with the same setting we have just described above.
@@
Figure 6 illustrates the topology of the problem and provides information regarding the maximal $H_{2}$-norm we computed by varying the parameter of the system. The node colors illustrate this latter aspect. In particular, darker colors correspond to the agents where the maximal $H_{2}$-norm turns out to be larger. It can be seen that certain groups of agents result in much larger $H_{2}$-norms. Therefore, they are much more important for the considered multi-agent system, as altering those agents may lead to an (almost) unstable system. Thanks to our novel solvers, this kind of analysis can now be accurately carried out at acceptable cost.
@@
In this example, we still consider agent dynamics defined by $2\times 2$ state matrices of dimension $m$. We aim at analyzing the parameter variation in off-diagonal elements of consecutive agents. The system matrices $A^{(i)}$, $B^{(i)}$, $C^{(i)}$, $K^{(i)}$, and $\mathbf{A}$ are defined as in Example 4.2 but here we consider $m=200$.
The Laplacian matrix $L$ of the underlying graph is given by Algorithm 2 and the corresponding matrices $B_{l}$ and $B_{r}$ determining the low-rank perturbation are defined in the following way. For a given odd index $31\leqslant k\leqslant 2m-$ 3, $B_{l},B_{r}\in{\mathbb{R}}^{2m\times 4}$ are
@@ 16       $L(i,i)=-\sum_{j=1}^{m}(L(i,j))$;
$\mathbf{A}^{(k)}(\mathsf{v})=\mathbf{A}-B_{l}\mathop{\mathrm{diag}}(v_{1},v_{1},v_{2},v_{2})B_{r}^{\mkern-1.5mu\mathsf{T}}.$
@@
We will study the behavior of the $H_{2}$-norm of the underlying system by varying the parameter $k\in\{61,141,221,301\}$. The total number of parameter configurations $\mathsf{v}=(v_{1},v_{2})$ for which $\mathbf{A}^{k}(\mathsf{v})$ is stable is equal to 6 04460446\,0446 044555As before, this number is obtained by multiplying all the possible values of $v_{1}$ and $v_{2}$, namely $40^{2}$, by all the adopted selections of $k$, i.e., $4$, and then subtracting the number of unstable cases (356).. Moreover, in this example, we considered the matrix $\mathbf{E}=I_{2m}$ for all cases, meaning that we measured the $H_{2}$-norm of the corresponding systems when all agents were independently externally disturbed.
@@
From the results in Table 3 we can see that our SMW+recycling Krylov scheme always attains very small relative errors. On the other hand, it turns out to be not very competitive in terms of computational timings. A finer tuning of the GCRO-DR parameters and the preconditioning operator may lead to some improvements in the performance of our scheme. Our projection framework outperforms all other approaches in terms of computational timing.
@@
In Figure 7 we display the magnitude of the $H_{2}$-norm by varying $\mathsf{v}$ for a fixed $k$. It can be seen how certain parameter values have a much greater impact on the $H_{2}$-norm of the system, thus emphasizing the important role of the corresponding pairs of agents. Once again, thanks to our novel solution processes, such analysis can be carried out in very few seconds on a standard laptop.
@@ 5 Conclusions
We proposed two different, efficient, and accurate methods for solving sequences of parametrized Lyapunov equations. The SMW+recycling Krylov approach is well-suited for small dimensional problems and is able to provide solutions achieving a very small relative errors. The proposed projection framework relies on the extended Krylov subspace method, and it makes use of the aforementioned SMW+recycling Krylov technique to solve the projected equations. Even though this second algorithm is tailored to large-scale problems, our numerical results show that it is extremely competitive also for medium-sized problems, especially if the number of Lyapunov equations to be solved is very large as it happens in the application settings we studied. We showed that the projection framework is able to speed up the entire solution process by an order of magnitude. Expensive analyses like viscosity optimization for vibrational systems and output synchronization of multi-agent systems are, thus, now affordable.
@@ Acknowledgements
The first author is a member of the INdAM Research Group GNCS. His work was partially supported by the funded project GNCS2023 “Metodi avanzati per la risoluzione di PDEs su griglie strutturate, e non” (CUP_E53C22001930001).